# ProS: Prompting-to-simulate Generalized knowledge for Universal Cross-Domain Retrieval

### 19 Dec 2023 v1

@@ 3.3 Retrieval by ProS
After two stage training, now we can generate CaDP by CaPS, and extract image features by CLIP with the help of CaPS.Fig. 2 (right) shows how to employ our proposed ProS for retrieval. Since we only need image features for retrieval, CLIP text encoder $g$ is discarded. We first use CaPS $\mathcal{M}$ to generate two CaDP, and then use CLIP Image Encoder $f$ to get well generalized image features $\mathbf{y}$:
@@
To perform retrieval, all images in $\mathcal{C}_{\text{test}}$ are indexed by Eq. 8 to formulate the gallery. Then, given a query, we do the same thing to obtain the query feature and produce a rank list by sorting similarities between query and all features in gallery. In the next section, we would give a comprehensive study about retrieval results of our proposed ProS.
@@ 4.1 Experimental setting
DomainNet [30] contains 596,006 images collected in 6 domains (Real, Sketch, Quickdraw, Infograph, Clipart, Painting) from 345 categories and split into 245, 55, 45 categories for training, validation and testing respectively. To satisfy the unseen domain requirement in both $\mathrm{U^{d}CDR}$ and $\mathrm{UCDR}$ protocol, leave-one-out evaluation protocol is applied, where we iteratively select one domain as unseen query set and use other domains for training. Additionally, for $\mathrm{U^{d}CDR}$ evaluation, we select 45 categories from training set and choose $25\%$ samples from unseen domains as queries (except for Quickdraw, which is $10\%$ due to it being too large). The gallery set is constructed with Real images from unseen categories or mixed with seen categories, termed Unseen Gallery and Mixed Gallery respectively.
Sketchy [35, 26] contains 75,471 Sketches and 73,002 images from 125 categories where the train, validation, and test splits contain 93 , 11 and 21 categories [37, 29], respectively. TU-Berlin [9, 26] has 20,000 Sketches and 204,489 images which are separated into 200 categories for training, 20 categories for validation, and 30 categories for testing [37, 29]. Both Sketchy and TU-Berlin are used for $\mathrm{U^{c}CDR}$.
Evaluation Metrics. For fair comparison, we utilize the same evaluation metrics following [29, 37]. For Sketchy and DomainNet, we evaluate precision and mean Average Precision w.r.t. top-200 candidates (Prec@200 and mAP@200). For TU-Berlin, we use Prec@100 and mAP@all instead as evaluation metrics.
Implementation Details. In all our experiments, we use a fixed image encoder and text encoder initialized with pre-trained CLIP (ViT-B/32). For Context-aware Prompt Simulator $\mathcal{M}$, we adopt a two layer ViT with random initialization. Following  [41], the length of text prompt $\mathbf{P_{t}}$ is set to 16 with 512 dimension. In addition, the dimension of other prompts including prompt units and Content-aware Dynamic Prompts are set to 768. The training epochs are set to 10 with early stopping of 2 epochs based on the evaluation performance for all the datasets with a batch size of 50. We use Adam optimizer with a learning rate of 1e-3 and adopt a cosine decay strategy for the learning rate.
@@ 4.2 Main Results
Baselines. We compare our ProS with two groups of methods, including two existing UCDR methods, i.e., SnMpNet [29] which adopt a ResNet as backbone and SASA [37], which implement with a ViT. For fair evaluation, we further build three CLIP-based baselines, including CLIP-full and two prompt tuning methods, i.e., CoOp [41] and VPT [17] where CLIP-full indicate fine-tunes full CLIP.
Results on UCDR and $\mathbf{U^{d}CDR}$. We firstly evaluate our ProS methods on DomainNet under UCDR and $\mathrm{U^{d}}\mathrm{CDR}$ settings. The UCDR results in Tab. 1 shows that our ProS method outperforms all baselines across all domains, demonstrating the superiority of ProS in terms of generalization on unseen domains and categories. Furthermore, we have highlight the following conclusions: First (CLIP vs. ViT): all CLIP-based methods outperform ViT-based method which prove our assumption that general knowledge gained in large-scale pre-training can enhance the model’s ability to handle the UCDR task. Specifically, CLIP-Full achieves 9.79% mAP@200 improvement than SASA. Second (Prompt Tuning vs. Fine-Tuning): VPT and CoOp all both outperforms CLIP-Full, demonstrating the effectiveness of prompt tuning to adapt CLIP. Third (ProS vs. Prompt Tuning): compared with VPT, our ProS further improves the performance by 3.21% mAP@200 on average, indicating that ProS effectively apply prompt tuning for UCDR task.
@@
Results on $\mathbf{U^{c}CDR}$. In Tab. 2, we explore the ability of ProS in handling semantic shift under $\mathrm{U^{c}CDR}$ setting on Sketchy and TU-Berlin. We observe that: First, our ProS shows consistent improvement compared to CLIP-Full and VPT, revealing that our ProS can enhance the ability of CLIP to manage semantic shifts and generalize to unseen categories. Second, compared to SASA, our ProS has a larger improvement on TU-Berlin than that on Sketchy($19.6\%$ vs. $0.81\%$ in mAP@200). We could see CLIP, CoOp and VPT, all of which use CLIP as the backbone, exhibit relatively low mAP on Sketchy, potentially caused by a poor performance of the visual backbone when handling this dataset. However, our method still surpasses all competitor and achieves the best performance.
@@ 4.3 Ablation Study
Impact of Each Components. To validate the effectiveness of each component in our ProS, we individually remove each component and experiment with the following five settings: 1) “w/o $SP$” (remove Semantic Prompt Units). 2) “w/o” $DP$ (remove Domain Prompt Units). 3) “w/o Mask” (remove mask operation in training). 4) “w/o $\mathcal{M}$” (remove Content-aware Prompt Simulator). 5) “w/o CLS” (remove CLS token). The experiment results are shown in Tab. 3. We observe that: First, removing either $SP$ or $DP$ led to a performance degradation of -2.65% and -2.56% in mAP@200 under mixed gallery settings, indicating their essential roles in Content-aware Prompt Simulation stage. Second, without the mask operation, mAP@200 has decreased by 1.27% which proves the effectiveness of mask operation. Third, when we remove CaPS $\mathcal{M}$ from our model, it lead to worse performance than full ProS, highlighting the significant benefits of $\mathcal{M}$ for UCDR task. Forth, training $CLS$ token is absolutely necessary as its omission caused a performance reduction of 6.18
@@
Table 3: Different component evaluated on DomainNet under UCDR protocol with Infograph as the unseen domain for queries.
Impact of Transformer Layer in $\mathcal{M}$. We further investigate the impact of Transformer layers in Context-aware Prompt Simulator $\mathcal{M}$. From the Tab. 4, it can be found that: 2 layer ViT yield the best performance. Although employing only 1 layer can slightly reduce computational costs, it results in a worse performance. Hence, for an optimal trade-off between accuracy and efficiency, a 2 layer ViT is chosen to implement $\mathcal{M}$.
@@
Table 4: Ablation study on effectiveness of different transformer layer in Context-aware Prompt Simulator $\mathcal{M}$.
Figure 3: Evaluation results of two prompts length. (a) investigate the impact of text prompt length. (b) analyse Content-aware Dynamic Prompt length generated by CaPS $\mathcal{M}$,where 0-0 represent VPT and 1-1 means one CaDP for domain and one for semantic.
@@
Figure 5: Retrieval Results under UCDR protocols on DomainNet. (a) displays the retrieval results of “Peas” by the query from a unseen query domain. (b) shows the retrieval results of a few queries from Quickdraw. True positives and false positives are shown with green and red borders, respectively.
Impact of two prompts length. We further conduct experiments to analyse the impact of two key prompt length in ProS. i.e.,text prompts and Content-aware Dyanmic Prompt (CaDP). Fig. 3(a) shows that when text prompts length increases, the performance improves to its peak when length is 16 and then gradually decreases. Therefore, we set the text prompt length to 16 for all datasets. In Fig. 3(b), we find that ProS achieve the state-of-the-art with only 1-1 CaDP. We speculate that this might be consistent with Prompt Units, which also use a single prompt to represent one domain or category.
@@ 4.4 Qualitative Analysis
To further demonstrate the effectiveness of our ProS, we visualize a few retrieval results produced by ProS. Fig. 5 (a) shows top five retrieved candidates of peas with query images from different domains, while Fig. 5 (b) shows top ten retrieved candidates with queries from Quickdraw. From the figure we can see that Sketch and Quickdraw images lack enough information such as color, texture, and details while Inforgraph has a object co-occurrence problem. They all obtain worse performance compared with other domains. This phenomenon will be exacerbated when the gallery contains both seen and unseen classes due to the interference caused by seen classes. Towards this phenomenon, we can conclude that: due to the quality issues of the dataset, the performance of our APL in certain domains may be relatively poor.
@@ 5 Conclusion
In this paper, to the best of our knowledge, we take the first attempt to deploy CLIP for Universal Cross-Domain Retrieval (UCDR) with a prompt tuning paradigm and propose a more generalized prompt method named Prompt-to-Simulate (ProS). Specifically,ProS can dynamically fits unknown domain and category distribution with guidance of source knowledge by a two-stage training paradigms following a mask-and-align objective. With the above approach, our model obtains quite strong performance under UCDR compared to state-of-the-art.

### 29 Feb 2024 v3

@@ 3.3 Retrieval by ProS
After two-stage training, now we can generate CaDP by CaPS, and extract image features by CLIP with the help of CaPS.Fig. 2 (right) shows how to employ our proposed ProS for retrieval. Since we only need image features for retrieval, CLIP text encoder $g$ is discarded. We first use CaPS $\mathcal{M}$ to generate two CaDP, and then use CLIP Image Encoder $f$ to get well-generalized image features $\mathbf{y}$:
@@
To perform retrieval, all images in $\mathcal{C}_{\text{test}}$ are indexed by Eq. 8 to formulate the gallery. Then, given a query, we do the same thing to obtain the query feature and produce a rank list by sorting similarities between the query and all features in the gallery. In the next section, we will give a comprehensive study of the retrieval results of our proposed ProS.
@@ 4.1 Experimental setting
DomainNet [30] contains 596,006 images collected in 6 domains (Real, Sketch, Quickdraw, Infograph, Clipart, Painting) from 345 categories and split into 245, 55, 45 categories for training, validation, and testing respectively. To satisfy the unseen domain requirement in both $\mathrm{U^{d}CDR}$ and $\mathrm{UCDR}$ protocol, leave-one-out evaluation protocol is applied, where we iteratively select one domain as unseen query set and use other domains for training. Additionally, for $\mathrm{U^{d}CDR}$ evaluation, we select 45 categories from training set and choose $25\%$ samples from unseen domains as queries (except for Quickdraw, which is $10\%$ due to its large size). The gallery set is constructed with Real images from unseen categories or mixed with seen categories, termed Unseen Gallery and Mixed Gallery respectively.
Sketchy [35, 26] contains 75,471 Sketches and 73,002 images from 125 categories where the train, validation, and test splits contain 93, 11, and 21 categories [37, 29], respectively. TU-Berlin [9, 26] has 20,000 Sketches and 204,489 images which are separated into 200 categories for training, 20 categories for validation, and 30 categories for testing [37, 29]. Both Sketchy and TU-Berlin are used for $\mathrm{U^{c}CDR}$.
Evaluation Metrics. For a fair comparison, we utilize the same evaluation metrics following [29, 37]. For Sketchy and DomainNet, we evaluate precision and mean Average Precision w.r.t. top-200 candidates (Prec@200 and mAP@200). For TU-Berlin, we use Prec@100 and mAP@all as evaluation metrics instead.
Implementation Details. In all our experiments, we use a fixed image encoder and text encoder initialized with pre-trained CLIP (ViT-B/32). For Context-aware Prompt Simulator $\mathcal{M}$, we adopt a two-layer ViT with random initialization. Following  [41], the length of text prompt $\mathbf{P_{t}}$ is set to 16 with 512 dimensions. In addition, the dimensions of other prompts including prompt units and Content-aware Dynamic Prompts are set to 768. The training epochs are set to 10 with early stopping of 2 epochs based on the evaluation performance for all the datasets with a batch size of 50. We use Adam optimizer with a learning rate of 1e-3 and adopt a cosine decay strategy for the learning rate.
@@ 4.2 Main Results
Baselines. We compare our ProS with two groups of methods, including two existing UCDR methods, i.e., SnMpNet [29] which adopt a ResNet as the backbone and SASA [37], which implement with a ViT. For fair evaluation, we further build three CLIP-based baselines, including CLIP-full and two prompt tuning methods, i.e., CoOp [41] and VPT [17] where CLIP-full indicates fine-tuning the full CLIP.
Results on UCDR and $\mathbf{U^{d}CDR}$. We first evaluate our ProS methods on DomainNet under UCDR and $\mathrm{U^{d}}\mathrm{CDR}$ settings. The UCDR results in Tab. 1 show that our ProS method outperforms all baselines across all domains, demonstrating the superiority of ProS in terms of generalization on unseen domains and categories. Furthermore, we have highlighted the following conclusions: First (CLIP vs. ViT): all CLIP-based methods outperform ViT-based method which proves our assumption that general knowledge gained in large-scale pre-training can enhance the model’s ability to handle the UCDR task. Specifically, CLIP-Full achieves 9.79% mAP@200 improvement than SASA. Second (Prompt Tuning vs. Fine-Tuning): VPT and CoOp both outperform CLIP-Full, demonstrating the effectiveness of prompt tuning to adapt CLIP. Third (ProS vs. Prompt Tuning): compared with VPT, our ProS further improves the performance by 3.21% mAP@200 on average, indicating that ProS effectively applies prompt tuning for UCDR.
@@
Results on $\mathbf{U^{c}CDR}$. In Tab. 2, we explore the ability of ProS in handling semantic shift under $\mathrm{U^{c}CDR}$ setting on Sketchy and TU-Berlin. We observe that: First, our ProS shows consistent improvement compared to CLIP-Full and VPT, revealing that our ProS can enhance the ability of CLIP to manage semantic shifts and generalize to unseen categories. Second, compared to SASA, our ProS has a larger improvement on TU-Berlin than that on Sketchy($19.6\%$ vs. $0.81\%$ in mAP@200). We could see CLIP, CoOp, and VPT, all of which use CLIP as the backbone, exhibit relatively low mAP on Sketchy, potentially caused by poor performance of the visual backbone when handling this dataset. However, our method still surpasses all competitors and achieves the best performance.
@@ 4.3 Ablation Study
Impact of Each Components. To validate the effectiveness of each component in our ProS, we individually remove each component and experiment with the following five settings: 1) “w/o $SP$” (remove Semantic Prompt Units). 2) “w/o” $DP$ (remove Domain Prompt Units). 3) “w/o Mask” (remove mask operation in training). 4) “w/o $\mathcal{M}$” (remove Content-aware Prompt Simulator). 5) “w/o CLS” (remove CLS token). The experiment results are shown in Tab. 3. We observe that: First, removing either $SP$ or $DP$ led to a performance degradation of -2.65% and -2.56% in mAP@200 under mixed gallery settings, indicating their essential roles in Content-aware Prompt Simulation stage. Second, without the mask operation, mAP@200 has decreased by 1.27% which proves the effectiveness of mask operation. Third, when we remove CaPS $\mathcal{M}$ from our model, it leads to worse performance than full ProS, highlighting the significant benefits of $\mathcal{M}$ for the UCDR task. Forth, training $CLS$ token is absolutely necessary as its omission caused a performance reduction of 6.18
@@
Table 3: Different components evaluated on DomainNet under UCDR protocol with Infograph as the unseen domain for queries.
Impact of Transformer Layer in $\mathcal{M}$. We further investigate the impact of Transformer layers in Context-aware Prompt Simulator $\mathcal{M}$. From the Tab. 4, it can be found that: 2 layer ViT yields the best performance. Although employing only 1 layer can slightly reduce computational costs, it results in a worse performance. Hence, for an optimal trade-off between accuracy and efficiency, a 2-layer ViT is chosen to implement $\mathcal{M}$.
@@
Table 4: Ablation study on the effectiveness of different transformer layers in Context-aware Prompt Simulator $\mathcal{M}$.
Figure 3: Evaluation results of two prompts length. (a) investigate the impact of text prompt length. (b) analyze Content-aware Dynamic Prompt length generated by CaPS $\mathcal{M}$, where 0-0 represents VPT and 1-1 means one CaDP for domain and one for semantic.
@@
Figure 5: Retrieval Results under UCDR protocols on DomainNet. (a) displays the retrieval results of “Peas” by the query from an unseen query domain. (b) shows the retrieval results of a few queries from Quickdraw. True positives and false positives are shown with green and red borders, respectively.
Impact of two prompts length. We further conduct experiments to analyze the impact of two key prompt lengths in ProS. i.e., text prompts and Content-aware Dynamic Prompt (CaDP). Fig. 3(a) shows that when text prompts length increases, the performance improves to its peak when the length is 16 and then gradually decreases. Therefore, we set the text prompt length to 16 for all datasets. In Fig. 3(b), we find that ProS achieves the state-of-the-art with only 1-1 CaDP. We speculate that this might be consistent with Prompt Units, which also use a single prompt to represent one domain or category.
@@ 4.4 Qualitative Analysis
To further demonstrate the effectiveness of our ProS, we visualize a few retrieval results produced by ProS. Fig. 5 (a) shows the top five retrieved candidates of peas with query images from different domains, while Fig. 5 (b) shows the top ten retrieved candidates with queries from Quickdraw. From the figure, we can see that Sketch and Quickdraw images lack enough information such as color, texture, and details while Inforgraph has an object co-occurrence problem. They all obtain worse performance compared with other domains. This phenomenon will be exacerbated when the gallery contains both seen and unseen classes due to the interference caused by seen classes. Towards this phenomenon, we can conclude that: due to the quality issues of the dataset, the performance of our APL in certain domains may be relatively poor.
@@ 5 Conclusion
In this paper, to the best of our knowledge, we take the first attempt to deploy CLIP for Universal Cross-Domain Retrieval (UCDR) with a prompt tuning paradigm and propose a more generalized prompt method named Prompt-to-Simulate (ProS). Specifically, ProS can dynamically fit unknown domain and category distribution with guidance of source knowledge by a two-stage training paradigm following a mask-and-align objective. With the above approach, our model obtains quite strong performance under UCDR compared to the state-of-the-art.